Factorize algebraic expressions: 15xy+2y-3x²y


1
Expert's answer
2021-02-24T05:58:25-0500
"-3x^2y+15xy+2y=y(-3x^2+15x+2)"



"D=(15)^2-4(-3)(2)=249"


"x=\dfrac{-15\pm\sqrt{249}}{2(-3)}"

"x_1=\dfrac{15-\sqrt{249}}{6}, x_2=\dfrac{15+\sqrt{249}}{6}"

"-3x^2y+15xy+2y"

"=-3y(x-\dfrac{15-\sqrt{249}}{6})(x-\dfrac{15+\sqrt{249}}{6})"